# Assessing Contamination in Large Language Models: Introducing the LogProber method

### v1

@@ 3.1 Can LogProber discriminate between newly designed and old ’Q-A’ items ?
The Cognitive Reflection Text (CRT) Frederick (2005b) is one of most popular psychological questionnaires used to investigate human reasoning and typically associated to the demonstration of two modes of reasoning, the fast "system 2" and slow "system 1" Kahneman (2011). The test is widely used in the scientific literature and can be found even beyond in numerous journal articles, blogs and forums. Therefore it is very likely that LLMs trained on web data containing the CRT. To validate LogProber, we deployed it on "classical" items of the CRT (oldCRT) and to more recently developed ones (newCRT), which has been published only once and so recently so they should not be represented in the training corpus of the considered LLMs Yax et al. (2024). The analysis shows that the A and B parameters for the 7 CRT items lied in linearly separable sectors of the log(A)/log(B) plane (Figure 2). More specifically log(A) was significantly higher (oldCRT: -1.04 newCRT: 1.28 t(11)=6.29, p<0.001) and log(B) was significantly lower (oldCRT: 3.38 newCRT: 0.37 t(9)=5.83, p<0.001) in the oldCRT compared to the newCRT test.
Looking more in details, LogProber returns higher scores for the first 3 oldCRT items, which are the first having being published and more wide-spread in the literature Frederick (2005a); Toplak et al. (2013). To sum up, LogProber was able to differentiate oldCRT items (present in the training corpus) from newCRT items (absent in the training corpus). These first results therefore provide first observational evidence of the effectiveness of LogProber.
@@ 3.2 Can LogProber discriminate between newly designed ’Q-A’ items before and after contamination ?
In the previous analysis, we showed promising evidence that LogProber is capable of discriminating between contaminated items and some conceptually equivalent, non-contaminated, versions (observational experiment). In the present section, we aim to provide experimental support to our claim by running dedicated fine-tuning experiments. To do so, we finetuned Llama 7B using both the questions and its answers (’Q-A’ tuning, see Table 2) the newCRT items, which corresponds to contaminating the model with them.
Results are reported in Figure 3 and show that the log(A) and log(B) parameter of LogProber lie in very different sector of the the plane, before and after fine-tuning.
The fine-tuning lead to high contamination log scores in the contaminated experiment (QA) as well as very high accuracies. On the other hand, in the reference experiment scores are very low and accuracy is average. This experiment validates the fact that contamination scores capture heavy model contamination if the model is contaminated in a Q-A manner. More specifically log(A) was significantly higher (STD=2.09, QA=-1.38, t(9)=5.19, p<0.001) and log(B) was significantly lower (STD=-0.43, QA=3.50, t(11)=4.45, p<0.001) in the fine-tuned compared to the native LLM test.
These results show that in a carefully controlled experiment, where we explicitely manipulated the training of a model to include the newCRT items (Q-A), LogProber was capable to detect that the model was finetuned with the newCRT Q-A items.
@@
Our first finetuning experiment featured a ’Q-A’ type of training, where the model was forced to learn to predict both the question and the answer of each individual CRT item. However, this represents only one way in which a model can be contaminated (See Table 2). For instance, the model can be trained on predicting the answer, based on the question (without further training the question per se). Such ’-A’ training strategy is very common to fine-tune the model on instructions Zheng et al. (2023); Taori et al. (2023). Conversely, a model could be trained only to predict the question, without necessarily including the (correct) answer in the string. Such ’Q-’ training strategy style would correspond to the presence in the corpus of a given question without the corresponding answer, for example when it is quoted to explain it or in research papers.
@@
This series of experiments show the complex relationship between markers of contamination (as derived from LogProber algorithm) and accuracy. On one side, (and somehow unsurprisingly) the accuracy is high whenever the model is trained to predict ’A’ given ’Q’, regardless of whether or not the ’Q’ itself is contained in the training. A language model can know by heart the answer to a question without knowing the question itself and can even show signs of surprise when reading the question. On the other side, LogProber detects contamination whenever the model has been trained to predict ’Q’, regardless of whether or not the ’A’ was included.
@@ 4 Discussion
In this paper we propose and validate a very simple algorithm, LogProber, which is capable to spot common forms of contamination in LLMs. More specifically, LogProber can be applied to detect contamination of text in a quesiton-answer (Q-A) format, which is commonly used both in benchmarks used in machine learning, as well as in psychology tests Srivastava et al. (2023); Yax et al. (2024); Coda-Forno et al. (2024).
LogProber analyses the (log)probability of the question in a Q-A sequence and is computationally cheap because, compared to other approaches, LogProber only requires a single forward pass on the question and storing the log-probabilities of tokens. As comparison, Dong et al. (2024) requires several generations of the answer (but doesn’t require access to the log-probabilies).
We demonstrated the effectiveness of LogProber by leveraging both observational and experimental evidence. Concerning observational evidence, we demonstrated the capacity of LogProber to differentiate between the classic items of a popular cognitive experiment (the Cognitive Reflection Test; Frederick (2005a); Toplak et al. (2013)) and newly designed versions of the same items Yax et al. (2024). More specifically, the contamination scores of the new items (in the form of log-parameters of an exponential function fitted in the log probabilities) were significantly different from those of the old items, which were, in turn, much better predicted by the LLMs. Concerning experimental evidence, we ran dedicated experiments where we trained (fine-tuned) a model to predict ’Q-A’ sequences of the newly designed items of the Cognitive Reflection Test. These experiments showed that LogProber was able to detect the changes in LLM training (i.e., it was capable of detecting its contamination).
For the sake of completeness, we ran other experiments where the additional only concerned either learning to predict only the question (’Q-’ training) or only the answer (’-A’ training). Critically, and somehow unsurprisingly, our results showed that LogProber was able to detect contaminations only when the Question was included in the training.
This result suggests that our approach is very effective in detecting contamination at the level of pre-training, since it is likely to occur in the ’Q-A’ format, and not in fine-tuning (likely to occur in the ’-A’ format: answer only) Wang et al. (2022); Taori et al. (2023); Zheng et al. (2023). This limitation does not substantially reduce the utility of LogProber, since in the open access LLM field pretraining dataset are rarely shared or very large to search a query in them. LogProber makes it possible to account for pretraining contamination at a very cheap cost.
When training a model using a full language modeling method - a common approach for pretraining autoregressive decoder-only LLM Wang et al. (2022) — the training data can follow three scenarios:
1.
The model is trained only on the question (Q) if the data does not include the answer.
2.
The model is trained on both the question and the answer (QA) if they appear consecutively in the data.
3.
Neither the question nor the answer is included in the training data (STD).
In this context, the QA scenario indicates data contamination. A low contamination score implies that neither Q nor QA scenarios are present in the training data, suggesting the STD scenario. Therefore, if a query results in a low contamination score, it is likely safe for use. Conversely, a high contamination score indicates that the model has been exposed to the question, suggesting potential contamination and that the model might be unreliable for that query.
In the open access community finetuning datasets are often shared and a lot smaller making it possible to know explicitly if the model was trained on the given query.
In short, LogProber can efficiently account for contamination acquired during pretraining (assuming the model is trained in a full language modelling - which is nearly always the case for modern LLMs) but may fail to capture finetuning contamination. This issue is mitigated by the fact that open access LLM finetuning datasets are often shared and fairly small making it possible to known explicitly if the model was trained on the given query.
Combining LogProber and Dong et al. (2024) might improve the results as it would make it possible to better differentiate A from STD and Q from QA but it still suffers from this disentangling between contamination and confidence.
Incidentally, these additional experiments also illustrate a very important, but often overlooked, fact: LLMs can be (over) trained to respond to a question, making them very familiar with answering the question without necessarily showing signs of knowing the question itself. Indeed, in our experiment ’-A’ training resulted in no sign of contamination, while the accuracy on the questions was extremely high. This also shows that LLM cognition is very different from human cognition especially in terms of how these two systems learn. Indeed, we can choose tokens on which LLMs learn while it is not possible to constraint the human learning mechanisms on specific words only.
LogProber presents other limitations, for instance, the algorithm returns a scores that can be difficult to interpret per se. To be interpreted they have to be compared with other sequences to put it in perspective and decide how likely the model is contaminated on the sequence. In our paper, we focus on an unambiguous case, where newly designed were compared to structurally equivalent old CRT, which served as an obvious reference point. This method (designing new items) could not be easily scalable to large benchmarks. However, by comparing LogProber scores across similar benchmarks, it will nonetheless be possible to make hypotheses about the relative likelihood of contamination. Furthermore, future experiments featuring more questions and LLMs could provide in future insights concerning the mapping between the LogProber scores and the likelihood of contamination.
This work also puts into perspective inherent limitations of quantifying contamination in LLMs in a black box setting on a given short sequence. Focusing only on the answer leads to mixing confidence and contamination while using the question solve this issue but only works in specific training settings. Further research could focus on doing both at the same time in order to increase the probability to get it right but we believe these methods would still not be foolproof and might require additional information such as neuronal activations from the network to accurately classify contaminated items.
@@
Another limitation of our approach is that it requires having access to the log-probabilities for the tokens and the question of the Q-A pair (typically used in benchmark and cognitive tasks). It could be, in principle, extended to cases where this information is not present by sampling each token in a sequence and deriving an "empirical" distribution. We note however that in that way the method would become more computationally expensive.
In conclusion, we propose and validate a method to estimate the likelihood of contamination for speficic string of text, based on a very computationally cheap analysis of the log-probabilities of the sequence. Our approach, LogProber, is particularly suited to detect contamination that should occur at the level of pretraining and can be easily applied in a domain-general manner, as soon as the cognitive test of the benchmarks quetions are framed in the ’Q-A’ manner.
Appendix A Training details
@@
While the alpaca default training data only fit on the response (like the paper) we formatted the newCRT data in different manner for the 3 training methods. For the QA-Training we included both question and answer in the response field so that the model fits on both. For the A-Training we included the question in the instruction field and the answer in the response field. Finally for the Q-Training we only included the question in the response field without the answer so that the model only fits on the question tokens.
Hyperparameters are the default hyperparameters for Llama 7B taken from the stanford alpaca git repository with a learning rate of 2e-5 and 3 epochs.

### v2

@@ 3.1 Can LogProber discriminate between newly designed and old ’Q-A’ items ?
To validate LogProber, we first deployed it on "classical" items of the CRT (oldCRT) and to more recently developed ones (newCRT). NewCRT items have been published recently, thus they should not be represented in the training corpus of the considered LLM Yax et al. (2024a). The analysis shows that the Safe Score for the 7 CRT items remains in different parts of the score line (Figure 2 upper part). More specifically, the score was significantly lower (oldCRT: -0.40 newCRT: 2.64 t(10)=9.70, p<0.001) in the oldCRT compared to the newCRT test. To sum up, LogProber was able to differentiate oldCRT items (present in the training corpus) from newCRT items (absent in the training corpus).
@@ 3.2 Can LogProber discriminate between newly designed ’Q-A’ items before and after contamination ?
In the previous analysis, we showed promising evidence that LogProber is capable of discriminating between contaminated items and some conceptually equivalent, non-contaminated, versions (observational experiment). In the present section, we aim to provide experimental support to our claim by running dedicated fine-tuning experiments. To do so, we finetuned Llama 7B using both the questions and its answers (’Q-A’ tuning, see Table 2) on the newCRT items, which corresponds to contaminating the model with them.
Results are reported in Figure 2 and show that the Safe Score of LogProber lies in very different sectors, before and after fine-tuning. The fine-tuning leads to low Safe Scores in the contaminated experiment (QA) as well as very high accuracies. On the other hand, in the control experiment, scores are very low and accuracy is average. This experiment validates the fact that scores capture model contamination if the model is contaminated in a Q-A manner. More specifically, the Safe Score was significantly lower (STD=2.98, QA=-0.83, t(8)=9.57, p<0.001) in the contaminated version (QA) compared to the original LLM test (STD).
@@
Our first finetuning experiment featured a ’Q-A’ type of training, where the model was forced to learn to predict both the question and the answer of each individual CRT item. However, this represents only one way in which a model can be contaminated (See Table 2). For instance, the model can be trained on predicting the answer, given the question (without being trained to memorize the question itself). Such ’-A’ training strategy is very common to fine-tune the model on instructions for example Zheng et al. (2023); Taori et al. (2023). Conversely, a model could be trained only to predict the question, without necessarily including the (correct) answer in the string. Such ’Q-’ training strategy style would correspond to the presence in the corpus of a given question without the corresponding answer, for example when it is quoted to explain it or in research papers.
@@
This series of experiments shows the complex relationship between markers of contamination (as derived from LogProber algorithm) and accuracy. On one side, (and somehow unsurprisingly) the accuracy is high whenever the model is trained to predict ’A’ given ’Q’, regardless of whether or not the ’Q’ itself is contained in the training. A language model can know by heart the answer to a question without knowing the question itself and can even show signs of surprise when reading the question. On the other side, LogProber detects contamination whenever the model has been trained to predict ’Q’, regardless of whether or not the ’A’ was included.
3.4 Comparison with CDD
As we believe contamination can be difficult to disentangle from confidence only looking at the answer, LogProber switches focus to detecting familiarity with the question rather than on the answer only. While this design works very well on QA types of finetuning, we have just shown its limitations on Q- and -A types of finetuning. It is now natural to compare this method with CDD, that only uses the answer to see how these two paradigms of contamination detection behave in practice on classical machine learning benchmarks.
As such, we first used MMLU Hendrycks et al. (2021) a MCQ type benchmark where we contaminated Qwen2.5-32B-Instruct Qwen et al. (2025) on part of the test set either following a QA (contaminated) or STD (control) type of training using QLoRA Dettmers et al. (2023). Our prediction is that Qwen2.5, being a top-end model, should be very good at MMLU thus should show very high confidence on this task. We separated a part of the test set into two splits : split A is included in the training data, while split B is not. Results are shown in Figure 4 left side. As expected, while LogProber is able to differentiate split A from split B in QA, CDD considers all splits to be contaminated on MMLU in all finetuning conditions even if the model is not contaminated (STD condition). See Appendix D for detailed comparison metrics, training details and different learning rates.
Now let’s compare both algorithms in a different task where models should show less confidence in answering the question. We contaminated a model on part of the BigCodeBench dataset, which contains coding problems where LLMs need to write a function corresponding to the given instructions. Figure 4 right side shows the results on the coding task. In this situation, CDD performs much better than in MMLU but seems still less effective than LogProber. See Appendix D for the detailed comparison metrics.
Even if LogProber appears on top of CDD in these experiments, it is important to emphasize the fact that here we only limited ourselves to the QA finetuning scenario. CDD would greatly outperform LogProber on Q- or -A types of finetuning.
@@ 4 Discussion
In this paper, we propose and validate a simple algorithm, LogProber, which is capable of detecting common forms of contamination in LLMs while disentangling confidence from contamination. It does so by analyzing the log-probability of the question, rather than focusing on the answer. This makes LogProber highly efficient compared to other approaches, as it requires only a single forward pass on the question. By comparison, CDD Dong et al. (2024) requires multiple answer generations—although it is worth noting that CDD does not require access to log-probabilities.
We demonstrated the effectiveness of LogProber using both observational and experimental evidence. For the observational evidence, we show that LogProber can distinguish between the widely known items of a popular cognitive experiment—the Cognitive Reflection Test—and newly designed variants of those same items Frederick (2005); Toplak et al. (2013); Yax et al. (2024a) . More specifically, the contamination scores for the new items were significantly different from those of the original items. As for the experimental evidence, we conducted dedicated experiments in which we fine-tuned a model on the ’Q-A’ sequences of newCRT items. These experiments demonstrated that LogProber was able to successfully detect contamination in the fine-tuned model. We also conducted additional experiments involving different fine-tuning scenarios: training the model to predict only the question (‘Q-’ training) or only the answer (‘-A’ training). Critically—and somewhat unsurprisingly—our results showed that LogProber was able to detect contamination only when the question was included in the training data. Third, we evaluated the method on classical machine learning benchmarks (MMLU and BigCodeBench), demonstrating near-perfect contamination detection in our experiments, with accuracy exceeding 98% in our experiments. In the same setting, CDD was less effective, frequently misidentifying uncontaminated scenarios as contaminated. This is due to the fact that CDD relies on the probability of the generated response, which can inherently conflate response confidence with contamination.
LogProber avoids this issue by evaluating the probability on the question, rather than focusing solely on the answer. However, it is also important to note that question-based methods like LogProber are well-suited for detecting ’Q-A’ contamination—typically introduced during pretraining—but are less effective at identifying answer-only (’-A’) contamination, which is more likely to occur during fine-tuning Zheng et al. (2023).
To sum up, question-based methods (such as LogProber) and answer-based methods (such as CDD) offer complementary advantages and limitations. Question-based methods can detect contamination originating during pretraining, providing valuable insights into the contents of the training corpus. However, they are generally unable to detect contamination involving only the answer, which typically arises during fine-tuning. Conversely, answer-based techniques struggle to disentangle contamination from confidence. As a result, they are less reliable for models that are genuinely strong or highly confident—an increasingly common situation with top-tier models on classical benchmarks like MMLU.
An interesting direction for future work is to combine these two approaches, aiming to leverage the strengths of each. However, this is not as straightforward as it may seem. For an answer-based method to flag contamination, the model must either be trained on the full ’Q-A’ pair, trained on the answer alone (’-A’ training), or simply be confident in the answer. For a question-based method to flag contamination, the model must have been trained on either the full ’Q-A’ pair or the question alone (’Q-’ training). As a result of these limitations, when evaluating the performance of a LLM, it is not possible to attribute its accuracy with absolute certainty to either high confidence or contamination. However, different combinations of outcomes from the question-based and answer-based methods can help constrain our interpretation of the origins of the LLM’s performance.
If both methods return positive (i.e., suggest contamination), the model may have been trained on the full ’Q-A’ pair, or it may have seen the question (’Q-’ training) and happens to be confident in its answer. (We cannot be 100% certain without additional assumptions). If the question-based method is positive but the answer-based method is negative, it suggests the model was trained on the question alone. If the answer-based method is positive but the question-based method is negative, it likely indicates ’-A’ training—i.e., the model was exposed to the answer during fine-tuning without seeing the question or that the model is just confident in its answer. Finally, if both methods return negative, we can reasonably conclude that the model was not contaminated on the item.
In summary, while answer-based methods cannot rule out the possibility that a model is simply confident, and question-based methods cannot rule out the possibility that a model has been fine-tuned in a specific manner, combining both approaches allows for more educated guesses about the presence and type of contamination and the origin of its performance.
To conclude, this work makes three main contributions. First, we propose and validate a method for estimating the likelihood of contamination in short text sequences, based on a computationally efficient analysis of sequence log-probabilities. Our approach, LogProber, is particularly well-suited for detecting contamination likely to occur during pretraining (’Q-A’ training). Second, we clarify an important distinction between contamination and confidence, and show how different types of contamination (’Q-A’, ’Q-’, and ’-A’) interact with different detection methods (question-based and answer-based), each with varying effectiveness depending on the fine-tuning scenario. Third, we discuss the upsides and drawbacks of question-based and answer-based approaches of contamination detection methods and how taking into account training techniques and model behavior is vital to account for contamination detection. This work could pave the way for new ways of detecting and thinking about contamination in efficient and effective manners—something much needed given the ever faster-evolving landscape of LLMs and the resulting need to monitor the evolution of their performance Yax et al. (2024b).
5 Limitations
@@
A limitation that we did not discuss yet of our approach is that it requires having access to the log-probabilities for the tokens and the question of the Q-A pair (typically used in benchmark and cognitive tasks). It could be, in principle, extended to cases where this information is not present by sampling each token in a sequence and deriving an "empirical" distribution. We note however that in that way the method would become more computationally expensive.
We would also like to emphasize the fact that LogProber focuses on how easily the LLM can predict a question, which is common when the model is contaminated. However, LLMs might also predict questions through in-context learning. For example, take the 178178… question — while the start is hard to predict, once the pattern emerges, a LLM can predict the end, resulting in a flatter line on the graph. This can be an issue for sequences with strong patterns, but in practice, such questions still score higher in Safe Score than truly contaminated ones, as the graph plateaus more slowly.
Finally, LogProber presents other limitations, for instance, the algorithm returns a score that can be difficult to interpret per se. To be interpreted, they have to be compared with other sequences to put it in perspective and decide how likely the model is contaminated on the sequence. In our paper, we set the threshold between contamination and Safe to 1 as it seemed to work very well in almost all our experiments.
We also did not test LogProber in implicit contamination scenarios, where the LLM is trained on questions very close to the one tested. This could be done in future work in order to test question-focused contamination detection techniques on this type of contamination and compare the results with algorithms like CDD that perform even on implicit contamination only looking at the answers.
Lastly, while we took care of using a model that was not familiar with our items in our first two experiments with Llama 1 7B, in the 3 experiment we used Qwen2.5-32B-Instruct that was developed after the publication of MMLU. We thus cannot be sure the test set we used to measure the contamination of the model already contained sequences the model was familiar with. We can at least claim that they were probably not in the pretraining set as LogProber didn’t catch any of these as contaminated. However, CDD might have caught some of them if indeed the model was trained on the MMLU test set during -A style finetuning during its post-training.
6 Acknowledgements
This work was granted access to the HPC/IA resources of [IDRIS HPE Jean Zay A100] under the allocation 2024- [AD011013693R2] and [IDRIS HPE Jean Zay H100] under the allocation 2024-[A0171011996] and made by GENCI. SP is supported by the European Research Council under the European Union’s Horizon 2020 research and innovation program (ERC) (RaReMem: 101043804), the Agence National de la Recherche (CogFinAgent: ANR-21-CE23-0002-02; RELATIVE: ANR-21-CE37-0008-01; RANGE: ANR-21-CE28-0024-01; ANR-17-EURE-0017), the Alexander Von Humboldt foundation, the EUR Frontiers in Cognition, the Idex PSL and a Google unrestricted gift. PYO is supported by ANR AI individual chair ANR-19-CHIA-0004.
Appendix A Llama 1 7B finetuning details
@@
While the alpaca default training data only fit on the response we formatted the newCRT data in different manner for the 3 training methods. For the QA-Training we included both question and answer in the response field so that the model fits on both. For the A-Training we included the question in the instruction field and the answer in the response field. Finally for the Q-Training we only included the question in the response field without the answer so that the model only fits on the question tokens. All these finetuning were done by finetuning all the weights of the model.
Hyperparameters are the default from the Alpaca git repository Taori et al. (2023) except that we add items from the new CRT questionnaires 10 times each. The questionnaire contains 7 questions and the model is trained for 3 epochs. Therefore, the model will see each of the 7 question-answer pair 30 times among the 52 000 Alpaca examples during the training process. These were run on a single GPU A100 taking 5 to 8 hours depending on the training set for each run.
Appendix B Old CRT and new CRT questionnaires
The list of items referred to as old CRT and new CRT are borrowed from Yax et al. (2024a) where the authors used original items from Frederick (2005) which language models are likely to be familiar with and made new fresh ones after the training of these models (thus LLMs trained before that time cannot be contaminated on them). Items are shown in Tab 2.
| oldCRT items | newCRT items |
Appendix C LogProber algorithm pseudocode
The pseudo code for LogProber is provided in Algorithm 1.
Appendix D Qwen2.5-7B-Instruct finetuning details and comparison with CDD
To contaminate Qwen2.5-32B-Instruct on MMLU we selected 10,000 samples from the train set of cais/mmlu on the huggingface hub as well as 100 samples from the test set that we duplicated 100 times (split A). The split B is made of 100 other samples from the test set that are not included in the training set. Finetuning is performed using QLoRA Dettmers et al. (2023) instead of full finetuning, as we did in the other experiments, because this LLM is a lot larger, and this finetuning was performed using 1 H100 and lasted approximately 8 hours for each model. The pipeline and hyperparameters are the default from alpaca Taori et al. (2023) explained in Appendix A with 5 epochs instead of 3. We tested different learning rates. The main paper shows results for a learning rate of 5e-5. Figure 5 shows results for a learning rate of 5e-4 yielding very similar results (see Table 3 for metrics), and Figure 6 for a learning rate of 5e-6. In this last figure, the learning rate is very small which doesn’t change very much the way the LLM is generating the question text. As such, LogProber has a hard time noticing contamination (as it is very light on the question). However, CDD seems to maintain its accuracy on the coding benchmark. While this is an unexpected behavior we believe it could come from the fact that when predicting the question, for the first few tokens the LLM doesn’t have enough context to predict them accurately, this naturally increases the Safe Score of LogProber as the cumulative logprobability curse is less horizontal. However, the answer being generated after the question, the LLM has enough context to notice it knows the generation by heart which may lead it to retrieve answers more easily than questions.
| Algorithm | Dataset | Learning Rate | Accuracy | Precision | Recall | F1 Score |
| CDD | MMLU | 5e-4 | 0.50 | 0.50 | 1.0 | 0.67 |
| LogProber | MMLU | 5e-4 | 1.0 | 1.0 | 1.0 | 1.0 |
| CDD | MMLU | 5e-5 | 0.50 | 0.50 | 1.0 | 0.67 |
| LogProber | MMLU | 5e-5 | 0.98 | 1.0 | 0.95 | 0.97 |
| CDD | MMLU | 5e-6 | 0.5 | 0.5 | 1.0 | 0.67 |
| LogProber | MMLU | 5e-6 | 0.57 | 1.0 | 0.13 | 0.23 |
| CDD | CODE | 5e-4 | 0.63 | 0.58 | 0.95 | 0.72 |
| LogProber | CODE | 5e-4 | 1.0 | 1.0 | 1.0 | 1.0 |
| CDD | CODE | 5e-5 | 0.68 | 0.62 | 0.93 | 0.74 |
| LogProber | CODE | 5e-5 | 0.99 | 1.0 | 0.99 | 0.99 |
| CDD | CODE | 5e-6 | 0.76 | 0.68 | 0.97 | 0.80 |
| LogProber | CODE | 5e-6 | 0.5 | 1.0 | 0.01 | 0.02 |
A similar setup was made for the coding benchmark with the first file of MBXP Athiwaratkun et al. (2023) being used to extract 10,000 samples for the training set and then 100 samples were taken from the bigcode/bigcodebench dataset for the test set. Similarly, they were duplicated 100 times each for the training set for split A, and another 100 were used to compose split B (which are not added to the training data).
For the STD training in MMLU we did only include the training set without splits A nor B making 10,000 datapoints in total. These were extracted from the question column for MMLU and we appended the answer column right after. For prompting reasons, because the question finishes with the "Answer:" prompt we added a left parenthesis afterwards to force the LLM to complete with the answer right after without reasoning. As the answer column starts with a left parenthesis we removed it to avoid having two of them. For the QA finetuning we included split A making 20,000 datapoints in total. For MBXP we used the content column for the train set and cut completions to the 1000th token in case they exceeded this length. For the splits A and B we used the complete_problem and canonical_solution columns appended after each other.
In order to evaluate CDD we need to stop the generation at some point to define the limit of what is an answer to a given question. While some LLMs generate an end of text token by themselves it is not always the case and we noticed finetuned Qwen models did not always put one and started hallucinating after the answer. Thus, we added a prompt asking the model to stop after the answer, and, because it was not very effective in MMLU, we cut the answers to one character for this benchmark (the model generates the answer in the first token of the completion). For BigCodeBench we did not restrict the answers and let the model generate up to 100 new tokens.
CDD hyperparameters were taken from the original paper Dong et al. (2024): is 0.05, is 0.01 and 50 answers were generated for each question with a temperature of 1 in addition to the greedily decoded answer.